OpenStudy (anonymous):

a rectangular field is to be enclosed by a fence. an existing fence will form one side of the enclosure. the amount of new fence bought for the other 3 sides is 1200ft. what is the maximum area of the enclosed field?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

the existing fence length was not given/

OpenStudy (lgbasallote):

?

OpenStudy (anonymous):

i have this much

OpenStudy (anonymous):

nope :\

OpenStudy (lgbasallote):

i smell calculus... @hartnn take over please

OpenStudy (anonymous):

yea its precalc lol

hartnn (hartnn):

u have 2x+y= 1200 and you need to maximize area xy , right ?

OpenStudy (anonymous):

umm...i need to the know the max area of the field

hartnn (hartnn):

can u take derivatives ?

OpenStudy (anonymous):

the answer is apparently 180,000ft^2 but idk how to get that

OpenStudy (anonymous):

i dont know what those are lol sorry

hartnn (hartnn):

if you are not taught derivatives, then this is not calculus! which specific topic does this question belong ?

OpenStudy (anonymous):

idk its basic functions and stuff

OpenStudy (anonymous):

its not calc its precalc

hartnn (hartnn):

ok, let me try 2x+y=1200 f(x) = x (1200-2x) = -2x^2 +1200 x now this is quadratic function

hartnn (hartnn):

compare it with f(x)=ax^2+bx+c then the maximum value of x occurs at -b/2a

OpenStudy (anonymous):

wha... L(*.*L)

hartnn (hartnn):

so here it will occur at x=1200/4 = 300 now find y= 1200-2(300) then find area =xy = ......

hartnn (hartnn):

which actually comes out to be 180000 ft^2

OpenStudy (anonymous):

its not supposed to be this complicated....idk what she (the teacher) wants me to do with this question....i see youre getting the right answer but surely this isnt the way im supposed to get it

hartnn (hartnn):

lol! u want complicated way ? use derivatives... this is the easiest i can think of......

OpenStudy (anonymous):

one sec ill try graphing

OpenStudy (anonymous):

it gives me x=600 which would make the area 360,000 :S

hartnn (hartnn):

u graphed -2x^2 +1200 x right ?

OpenStudy (anonymous):

i dont understand how you made it into a quadratic, could u explain that?

hartnn (hartnn):

2x+y=1200 , u got right ?

OpenStudy (anonymous):

howd you make 2x+y=1200 into that?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so y=-2x+1200

OpenStudy (anonymous):

howd u get the square?

hartnn (hartnn):

AREA = length * breadth = x*y put that y in this

OpenStudy (anonymous):

Vertex of parabola, right?

OpenStudy (anonymous):

so wait we're comparing 2x+y=1200 and 2x*2y=A A-1200=y (2x*2y)-(2x+y)=y 2x2y-2x+y=y k now i just confused myself lol

hartnn (hartnn):

y=-2x+1200 f(x) = xy = x(-2x+1200) = ??

OpenStudy (anonymous):

2x+y=1200 y=1200-2x area = xy = 1200x-2x^2

OpenStudy (anonymous):

ooohh...okay so x*y=x(1200-2x)=-2x^2+1200x k ill graph that

hartnn (hartnn):

and yes, find vertex of that parabola

OpenStudy (anonymous):

x=-1200/-4 x=300 so that means y=600 and that gives 180,000 YAY T(^_^T)

hartnn (hartnn):

u got it :)

OpenStudy (anonymous):

thanks so much hartnn :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

^_^